OpenStudy (anonymous):

9. If Steven can mix 20 drinks in 5 minutes, Sue can mix 20 drinks in 10 minutes, and Jack can mix 20 drinks in 15 minutes, how much time will it take all 3 of them working together to mix the 20 drinks? A. 2 minutes and 44 seconds B. 2 minutes and 58 seconds C. 3 minutes and 10 seconds D. 3 minutes and 26 seconds E. 4 minutes and 15 seconds The answer is A but I don't understand why.

OpenStudy (sriram):

less than 3 mins ans i got is 30/11 mins

OpenStudy (lgbasallote):

\[\frac{t}{5} + \frac{t}{10} + \frac{t}{15} = 1\] solve for t :)

OpenStudy (lgbasallote):

you get the solution btw?

OpenStudy (sriram):

30/11=2.7272 is 2 mins 43 secs

OpenStudy (sriram):

so prob is d 1st opt

OpenStudy (anonymous):

Where does the 30 and 11 come from?

OpenStudy (sriram):

solve eq given above by lgbasallote

OpenStudy (lgbasallote):

multiply the equation i presented by 30 \[6t + 3t + 2t = 30\]

OpenStudy (lgbasallote):

if you add it up \[11t = 30\]

OpenStudy (lgbasallote):

satellite shall demonstrate to us his awesomeness now...

OpenStudy (anonymous):

Okay now I understand the equation and why the answer was given. Now it gives me 2.727272, how do I put that in minutes and seconds?

OpenStudy (phi):

first, how many seconds in a minute? if you have 0.7272 minutes, you have a fraction of a minute. to see how to change this to seconds, think about this problem 0.5 minutes (one half of a minute) is 30 seconds. notice that 0.5*60= 30

OpenStudy (anonymous):

Okay so far it's 30 seconds and since it's 2.7272, We have 2 minutes and 30 seconds. So if 0.5=30seconds then 0.25=15seconds?

OpenStudy (phi):

first, how many seconds in a minute?

OpenStudy (anonymous):

2 minutes and 30 seconds. 30+15=45 so it's 2 minutes and 45 seconds.

OpenStudy (anonymous):

60 seconds in 1 minute

OpenStudy (phi):

I wanted you to notice that 0.5 times 60 = 30 so using the same idea, 0.7272 times 60 will be the answer your idea that 0.5 minutes + 0.25 minutes is 0.75 minutes or 30+15= 45 seconds and 0 .75 is close to 0.7272 so 45 seconds is close to the actual answer

OpenStudy (anonymous):

Okay and it gave me 43.632 (0.7272 x 60) and then you round it which gives you 44. Okay thank you guys so much.

OpenStudy (phi):

yes.

OpenStudy (phi):

notice that if you have 2 minutes, you figure out how many seconds by multiplying 2* 60= 120 seconds if you have a fraction (decimal) like for example 0.7272, you do the same thing, multiply: 0.7272*60 = 43.632 seconds or 44 rounded.